1) The observed numbers for dominant and recessive types in an F2 generation are 154 and 46. You would assume the expected numbers to be in a ---------- to 1 ratio.

would the answer for question 1 be a 3 to 1 ratio i'm really confused.

2) Looking at the observed numbers in question 1 what would your expected number of dominant types be?

I don't know please help.

3) Using the same observed numbers in question 1 what would the expected number of recessive types be?

4) Using the observed numbers from question 1 what would the expected number of individuals with a dominant phenotype and a heterozygous genotype? Remember the expected genotypic ratio is 1 (AA): 2 (Aa): 1 (aa).

Please answer all the questions and if you can explain this is due today thank you so much for your help I really appreciate it .
Answer #1

1. Phenotypic ratio expected is 3 (dominant): 1 (recessive)

2. Total number of progeny is 154 + 46 = 200

  • Thus the genotypic ratio to expected is 3:1
  • Therefore expected number of dominant phenotype is = 150
  • As ratio is 3:1 therefore 3+1 = 4 and 4x = 200. The value of x = 50
  • Thus dominant phenotype is 3 × 50 = 150

3 . Number of recessive 1 × 50 = 50

4. Total number of Dominant phenotype = 156

  • This number of progeny has either AA or Aa genotype in ratio of 1:2 (because genotypic ratio should be 1:2:1)
  • Therefore 1+ 2 = 3 and 3y = 156 and therefore y = 52
  • Therefore number of Dominant phenotype is 1y = 52
  • Number of Heterozygous Genotype is 2y = 104